OpenStudy (anonymous):

Explain why the Distributive Property is sometimes called the Distributive Property of Multiplication Over Addition.

OpenStudy (anonymous):

it's because Distributive Property lets you multiply a sum by multiplying each addend separately and then add the products.

OpenStudy (anonymous):

The distributive property of multiplication over addition states that when the sum of two numbers is multiplied by a number, then the first number is distributed to those two numbers and it is multiplied by each of them separately. The distributive property states that multiplying the sum of two numbers by another number, is equivalent to multiplying each of the addends by that number and then adding the products together. The Distributive Property stands for the product of a number where the sum is equal to the sum of the individual products of the addends and the number. a * (b + c) = (a * b) + (a * c) 3 * (4 + 5) = (3 * 4) + (3 * 5)

OpenStudy (anonymous):

it's because Distributive Property lets you multiply a sum by multiplying each addend separately and then add the products. 4 x 53 (4 x 50) + (4 x 3) 200 + 12 212

OpenStudy (jiteshmeghwal9):

Becoz when any number is distributed btw any two numbers 2(4+6) 2(4)+2(6) 8+6 14

OpenStudy (anonymous):

How do you simplify this one.

OpenStudy (jiteshmeghwal9):

which one @CrazyC246 ????

OpenStudy (anonymous):